Suppose x is a normally distributed random variable with μ=48 and σ=5. Find a value x0 of the random variable x that satisfies the following equations or statements.

a.​ 10% of the values of x are less than x0.

b.​ 1% of the values of x are greater than x0.


1
Expert's answer
2021-05-07T08:53:26-0400

"\mu=48"

"\sigma=5"

Use the distribution function of the standard normal distribution

"\displaystyle\Phi(x)=\dfrac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^x e^{-\frac{t^2}{2}}dt"

Note that

"\Phi(-\infty)=0"

"\Phi(+\infty)=1"


a. "P(X<x_0)=0.1"


"P(-\infty<X<x_0)=0.1"

"\Phi\left(\dfrac{x_0-\mu}{\sigma}\right)-\Phi(-\infty)=0.1"


"\Phi\left(\dfrac{x_0-48}{5}\right)-0=0.1"


"\Phi\left(\dfrac{x_0-48}{5}\right)=0.1"


"\dfrac{x_0-48}{5}=\Phi^{-1}(0.1)"


"x_0-48=5\cdot \Phi^{-1}(0.1)"


"x_0=5\cdot \Phi^{-1}(0.1)+48"


"x_0\approx5\cdot (-1.28)+48=-6.41+48=41.59"



b. "P(X>x_0)=0.01"


"P(x_0<X<+\infty)=0.01"


"\Phi(+\infty)-\Phi\left(\dfrac{x_0-\mu}{\sigma}\right)=0.01"


"1-\Phi\left(\dfrac{x_0-48}{5}\right)=0.01"


"-\Phi\left(\dfrac{x_0-48}{5}\right)=0.01-1"


"\Phi\left(\dfrac{x_0-48}{5}\right)=1-0.01"


"\Phi\left(\dfrac{x_0-48}{5}\right)=0.99"


"\dfrac{x_0-48}{5}=\Phi^{-1}(0.99)"


"x_0-48=5\cdot\Phi^{-1}(0.99)"


"x_0=5\cdot\Phi^{-1}(0.99)+48"


"x_0\approx5\cdot2.33+48=11.63+48=59.63"